What is the percent composition of aconitine, C34H47NO11?



% C
% H
% N
% O
1
Expert's answer
2020-10-12T13:34:07-0400

Given molicular formula = C34H47NO11

In this compound the ratio of number of moles of constituent atoms 

     C : H : N : O = 34 : 47 : 1 : 11




So, number fraction of moles of C atoms = 34/(34+47+1+11)

       =( 34/93 )

 So, percent of composition of C atom = (34/93)×100

                       =36.559%

     (upto three nearest decimal place)




Similarly, number fraction of moles of H atoms = 47/(34+47+1+11)

       =( 47/93 )

        So, percent of composition of C atom = (47/93)×100

                       =50.538%

      (upto three nearest decimal place)




So, number fraction of moles of N atoms = 1/(34+47+1+11)

       =( 1/93 )

 So, percent of composition of C atom = (1/93)×100

                       =1.075%

     (upto three nearest decimal place)




So, number fraction of moles of O atoms = 11/(34+47+1+11)

       =( 11/93 )

 So, percent of composition of O atom = (11/93)×100

                       =11.828%

       (upto three nearest decimal place)




Hence, 

the percent composition of C34H47NO11 is......

     

     36.559%  C

      

     50.538%  H


     1.075%   N

  

    11.828%  O